# Forging Tokens for Improved Storage-efficient Training

### v1

@@ 4 Experiment
In this section, we validate the effectiveness of our proposed augmentation strategies across various scenarios where data augmentation is crucial. Initially, we evaluate our method’s performance in the context of storage-efficient ImageNet-1k classification. Furthermore, we demonstrate the adaptability of token-based vision model training to dense prediction tasks, showcasing additional improvements facilitated by our augmentation strategies. Additionally, we present experimental results illustrating how our strategies contribute to enhanced model generalizability in fine-grained classification and robustness evaluation. Lastly, we explore the applicability of our method to alternative input formats, demonstrating its extensibility. Unless otherwise stated, all the evaluations are performed on SeiT [36] framework.
4.1 Storage-efficient ImageNet
Table 2 summarizes the performance and compression ratio of storage-efficient ImageNet-1k classification. Our method is compared with image-based approaches, including uniform random sampling, C-score [26]-based sampling, adjusting image resolution, and adjusting JPEG quality factor. Despite a 1% compression ratio, our method achieves a 75.5% top-1 accuracy, outperforming various image-based approaches. In Table 3, our method consistently outperforms the token-based baseline by 1.5–4.1%p in top-1 accuracies. Notably, the effectiveness of our method becomes more pronounced as the available storage size decreases, emphasizing its efficiency in low-storage scenarios.
| Input |
Baseline | Ours | ||
| Token | 1.4 GB | 74.0 | 75.5 (+1.5) | ||
| 1.1 GB | 70.2 | 73.1 (+2.9) | |||
| 0.8 GB | 66.3 | 69.1 (+2.8) | |||
| 0.5 GB | 59.7 | 63.8 (+4.1) | |||
| 0.3 GB | 47.2 | 51.2 (+4.0) |
@@ 4.1 Storage-efficient ImageNet
| Baseline | 93.5 | 79.7 | 43.1 | 50.1 |
| Ours | 93.6 | 83.9 | 47.5 | 55.5 |
| + | +0.1 | +4.2 | +4.4 | +5.4 |
@@
| Baseline | 74.0 | 50.7 | 62.6 | 25.5 | 22.6 | 15.7 |
| Ours | 75.5 (+1.5) | 58.6 (+7.9) | 66.8 (+4.2) | 30.2 (+4.7) | 27.7 (+5.1) | 18.3 (+2.6) |
4.2 Fine-grained classification
To further investigate the generalizability of our method, we conduct experiments on four different fine-grained datasets: Flowers [35], StandfordCars [27], iNat-18 [50], and iNat-19 [51]. Specifically, we tokenize each dataset and then fine-tune the token-trained model (ViT-B) on ImageNet-1k with the tokenized dataset. Since fine-grained datasets require classifying objects with intricate characteristics, they are particularly sensitive to undesirable noise (i.e., RRC in SeiT) or excessive perturbation (i.e., Emb-Noise) resulting from data augmentation in token embedding space. Table 4 shows the quantitative performance (top-1 accuracies) on fine-grained datasets. Our approach consistently outperforms the performance of the baseline across various datasets. These results confirm the effectiveness of our proposed method in fine-grained classification.
@@ 4.3 Evaluation on robustness datasets
We further evaluate the robustness of our method on a set of benchmarks, such as adding Gaussian noise [22] or Gaussian blur [22], ImageNet-R [23], Sketch [52], and ObjectNet [5]. The results in Table 5 demonstrate that our method consistently outperforms the baseline in terms of robustness accuracies. Previous studies [7, 47] observe that the key to the input pixel robustness depends on the pixel-level augmentations. Since our method enables pixel-based data augmentation to tokens, it achieves higher robustness accuracies than the baseline, which does not effectively exploit the pixel-based data augmentation.
@@
| Baseline | 77.3 | 70.3 | 53.3 | 43.2 |
| Ours | 81.4 | 76.5 | 61.6 | 52.3 |
| + | +4.1 | +6.2 | +8.3 | +9.1 |
4.4 Classification with limited data
Data augmentation becomes more critical in limited data scenarios. To assess the impact of our approach in such situations, we conduct experiments on the ImageNet-100 benchmark by randomly sampling data based on varying data ratios. The results (top-1 accuracies), as depicted in Table 6, consistently demonstrate our method’s superiority over our baseline. Notably, the performance gap increases as the amount of training data decreases. With only 13k images, our method exhibits an improvement of approximately 21% over the baseline, indicating the effectiveness of our approach, especially in data-scarce scenarios.
4.5 ADE-20k semantic segmentation
A limitation of token-based learning is often substantiated by learning dense prediction tasks. We conjecture that our method could facilitate training for those tasks, such as segmentation tasks. We extract tokens from ADE-20k [62] and utilize the pre-trained ViT-B/16 model on tokenized ImageNet-1k to evaluate the effectiveness of our method in semantic segmentation. Subsequently, we fine-tune the pre-trained model using the tokenized ADE-20k dataset. Table 7 reports the mean intersection-over-union (mIoU) as the evaluation metric. The results demonstrate a notable performance increase (+2.5%p mIoU) with our data augmentation strategies compared to the baseline. These findings highlight the significance of addressing artifacts introduced by directly applying pixel-based data augmentation to token embeddings in semantic segmentation, a challenge effectively mitigated by our augmentation approach.
4.6 Analysis of TokenAdapt
We analyze the influence of our approach on training ViT-B for token-based ImageNet-1k classification, as depicted by the loss curves in Figure 6. Observing the training loss, our method consistently exhibits higher values than the baseline. Meanwhile, the test loss aligns with the trends of the baseline. This implies that the integration of token augmentation with our method effectively enhances data diversity, mitigating overfitting concerns. Notably, while the baseline’s test loss increases in later training epochs, our approach consistently decreases, highlighting its effectiveness in preventing the overfitting problem.
Furthermore, to demonstrate the efficacy of our strategies, we visualize the embedded feature representation in Figure 7. This illustration offers insight into the linear separability of augmented data points, particularly when pixel-based data augmentation is applied with our method. When pixel-based data augmentation is directly applied to token embedding without our method, there is significant overlap among augmented data points from different classes, indicating low linear separability. This overlap suggests that augmenting tokens without our method leads to substantial shifts in the training data distribution. Maintaining distributional similarity between clean and augmented data is crucial for model performance, as witnessed in Cubuk et al. [13], and significant distribution shifts can result in performance degradation. In contrast, our method shows high linear separability, demonstrating its effectiveness in diversifying the data while minimizing the distributional gap between clean and augmented data. Moreover, this behavior aligns with the observed trend in the case of applying data augmentation to images (DeiT-B [48] in Figure 7).
4.7 Ablation study
@@
Impact of each augmentation strategy. In Table 8, we report the impact of proposed token augmentations. Notably, integrating ColorAdapt results in a 1.0%p performance boost, emphasizing the positive effect of preserving object structure during model training. Additionally, incorporating TokenAdapt yields an extra 3.1%p improvement. These results confirm that exploring token augmentations is a key factor in fully exploiting pre-extracted tokens.
| Input | Baseline | Ours |
@@
Extensibility of our approach. In this section, we conjecture the broad applicability of our method across various scenarios. To explore its extensibility, we delve into two distinct situations. Firstly, we examine the integration of our method with an alternative tokenizer—the OpenImages-trained VQGAN tokenizer [17], alongside the ImageNet-1k-trained ViT-VQGAN tokenizer [56]. The results in Table 9 consistently demonstrate improved performance on ImageNet-100 across different tokenizers. Additionally, we investigate the adaptability of our approach to a different input format by focusing on Discrete Cosine Transform (DCT) coefficients. Inspired by JPEG compression, we apply DCT to each image patch, yielding quantized DCT coefficients utilzed as input data during model training. As shown in Table 9, our approach enhances top-1 accuracy, similar to the tokenizer scenario. This consistent improvement across diverse input formats underscores the universal applicability of our method, highlighting its potential as a widely applicable solution for diverse input representations.
@@ 4.8 Implementation details
We use the pre-trained ViT-VQGAN tokenizer [56] for token extraction. For our TokenAdapt module, we employ a single transformer block for conversion and reverse function. The TokenAdapt module is trained for one epoch on ImageNet-1k with a batch size of 128, a learning rate of 0.001 with cosine scheduling, and a weight decay of 0.05. We follow the training recipes from SeiT [36] and mmsegmentation [9] for token-based classification and semantic segmentation, respectively. Detailed hyperparameters and settings are available in the Appendix.
@@ 5 Conclusion
In this paper, we address a significant challenge in token-based vision model training: the limitations of integrating existing data augmentations designed for pixel images into token embedding due to input domain shift. To address this challenge, we propose simple yet effective token-based augmentation strategies, namely TokenAdapt and ColorAdapt. TokenAdapt realigns token embedding space to another embedding space compatible with pixel-based data augmentations. Once trained, the TokenAdapt module can be applied to different datasets and scenarios without the need for additional training. Additionally, we introduce ColorAdapt, a color-changing augmentation for tokens, to further diversify the training data. Our experimental results across diverse scenarios, including storage-efficient ImageNet-1k classification, fine-grained classification, robustness benchmarks, and ADE-20k semantic segmentation, consistently validate the effectiveness of our proposed method. Additionally, the consistent performance improvement across diverse input formats highlights the potential of our approach as a widely applicable solution for various input representations.
Appendix
@@
In this supplementary material, we describe more details of our method. In Section A, we include more implementation details for our experiments (i.e., token-based image classification and semantic segmentation). Section B explores the data efficiency (Section B.1) and computational cost (Section B.2) of our TokenAdapt module. Additionally, Section C presents more qualitative results showcasing the effectiveness of our proposed TokenAdapt and ColorAdapt for augmenting tokens.
Appendix A Implementation details
Token-based image classification. For token ImageNet-1k training, we follow the training recipe from SeiT [36], adjusting only the warm-up epochs for stable convergence. We conducted token-based image classification on the ViT-B/16 model [16, 48] and used a learning rate of 0.0015 with cosine scheduling and a weight decay of 0.1. The model was trained for 300 epochs with a batch size of 1024. Regarding data augmentations, SeiT [36] incorporates Token-RRC, Token-EDA, Token-CutMix, and Emb-Noise. As a default, we integrate these token-based data augmentation methods and further enhance token diversity by applying additional augmentations using our TokenAdapt and ColorAdapt. Specifically, we apply pixel-based data augmentations (e.g., RRC, hFlip, affine transformations, Mixup, and CutMix) to tokens using our TokenAdapt module with a probability of 0.5. We adopt the hyperparameters for data augmentation proposed in DeiT [48]. For token-based fine-grained classification, following the SeiT, we used the training recipe of DeiT. In all experiments, when the number of data points decreased, we modified the number of total training iterations to ensure a fair comparison.
Token-based semantic segmentation. For the token ADE-20k dataset preparation, we initially resize the entire ADE-20k dataset to . Following the procedure described in SeiT, we utilize the ViT-VQGAN tokenizer [56] to extract tokens from the resized images. Token-based semantic segmentation on ADE-20k follows the training recipe from mmsegmentation [9], using the DeiT-B/16 model with UperNet [54]. The training involves a learning rate of 6e-5 with polynomial scheduling and a weight decay of 0.01. The model was trained for 80k iterations with a batch size of 16. Regarding data augmentations, as mentioned in Section 3.2, we do not employ mixup-based augmentations in token-based semantic segmentation. Furthermore, we observe that token-level Random Resized Crop (RRC) does not enhance performance due to the importance of maintaining structural integrity in pixel-level classification tasks. Thus, we apply geometric pixel-based data augmentations only using our TokenAdapt module.
| # of training data | 100 | 1k | 5k | 256k | 1,281k |
@@ Appendix A Implementation details
Appendix B Efficiency analysis of TokenAdapt
B.1 Data efficiency of TokenAdapt
To investigate the data efficiency of the TokenAdapt module, we explore the impact of reduced training data for the TokenAdapt module training. Specifically, We vary the number of training data for the TokenAdapt module training and subsequently train a storage-efficient ImageNet-1k classification model using the trained TokenAdapt module. Table A shows that even with a reduced number of training data for our TokenAdapt module, the performance degradation is minimal. Notably, the top-1 accuracy remains more than 1%p higher than the baseline (74.0% top-1 accuracy) even with only 100 training data. This robust performance suggests that our TokenAdapt module is not significantly affected by variations in the training data size.
B.2 Computational costs
@@
To validate the computational efficiency of TokenAdapt, we compare ViT-S models trained on ImageNet-100 with different training strategies (Table B.2). Three scenarios are considered: (1) only token-based augmentations (baseline), (2) pixel-based augmentations in decoded images (baseline), and (3) pixel-based augmentations using our TokenAdapt module. For baseline, utilizing the ViT-VQGAN decoder (86M) and encoder (86M) during each forward computation significantly increases training time by over 8 times compared to the baseline, making it impractical for addressing the data augmentation challenge. In contrast, our TokenAdapt module, requiring only 3.7M parameters, increases training time by only 8% compared to the baseline while achieving a remarkable performance improvement over 3%p in top-1 accuracy. Notably, TokenAdapt even outperforms baseline, indicating that the full decoding-augmentation-encoding process may introduce undesirable noise during tokenization. These results demonstrate the efficient handling of the data augmentation challenge by our TokenAdapt module. We note that the TokenAdapt module can be removed during the inference phase.
Appendix C More qualitative examples

### v3

@@ 4 Experiment
In this section, we validate the effectiveness of SeiT++ on various scenarios. First, we evaluate the performance of our method in the context of storage-efficient ImageNet-1k classification and transfer learning on fine-grained datasets. Then, we demonstrate the adaptability of token-based vision model training to ADE-20k semantic segmentation, showing additional improvements facilitated by SeiT++. In addition, we present experimental results illustrating how our strategies contribute to enhanced model generalizability in robustness evaluation. We also explore the applicability of our augmentation strategies to alternative input formats, demonstrating its extensibility. Unless otherwise stated, SeiT++ refers to the combination of MTM and our token augmentation strategies.
| Method | Input | # of images | Top 1 Acc. |
| Full pixels | Image | 1,281 k | 81.8 | 140.0 GB | ||
| Uniform random sampling | Image | 512 k | 74.0 | 54.6 GB (39%) | ||
| C-score [28] based sampling | 512 k | 73.3 | 53.3 GB (38%) | |||
| Adjusting JPEG quality factor to 5 | 1,281 k | 74.6 | 11.0 GB (8%) | |||
| Adjusting image resolution ( 0.2) | 1,281 k | 75.2 | 9.6 GB (7%) | |||
| SeiT [38] | Token | 1,281 k | 74.0 | 1.4 GB (1%) | ||
| SeiT++ | 1,281 k | 77.8 | 1.4 GB (1%) |
4.1 Storage-efficient ImageNet-1k
Table 1 summarizes the performance and compression ratio of storage-efficient ImageNet-1k classification. Our method is compared with image-based approaches, including uniform random sampling, C-score [28]-based sampling, adjusting image resolution, and adjusting JPEG quality factor. Despite a 1% compression ratio, when a ViT-B/16 model is pre-trained on tokenized ImageNet-1k using MTM and the ViT model is fine-tuned, our method achieves a 77.8% top-1 accuracy, outperforming various image-based approaches. In Table 2, our method consistently outperforms the token-based learning baseline SeiT by 1.5–6.7%p in top-1 accuracies. Notably, the effectiveness of our method becomes more pronounced as the available storage size decreases, highlighting its efficiency in low-storage scenarios. Furthermore, SeiT++ shows more significant performance gains with MTM. For a storage size of 1.4GB, SeiT++ achieves an improvement of 1.5%p over SeiT without MTM, while with MTM the improvement increases to 2.7%p. These results confirm that SeiT++ effectively takes advantage of MTM.
| Storage | SeiT | SeiT++ | ||
| w/o MTM | w/ MTM | w/o MTM | w/ MTM | |
| 1.4 GB | 74.0 | 75.1 | 75.5 (+1.5) | 77.8 (+2.7) |
| 1.1 GB | 70.2 | 74.1 | 73.1 (+2.9) | 76.5 (+2.4) |
| 0.8 GB | 66.3 | 70.6 | 69.1 (+2.8) | 74.1 (+3.5) |
| 0.5 GB | 59.7 | 64.5 | 63.8 (+4.1) | 70.0 (+5.5) |
| 0.3 GB | 47.2 | 53.9 | 51.2 (+4.0) | 60.6 (+6.7) |
@@ 4.1 Storage-efficient ImageNet-1k
| SeiT [38] | 93.5 | 79.7 | 43.1 | 50.1 |
| SeiT++ | 95.2 (+1.7) | 86.9 (+7.2) | 52.4 (+9.3) | 58.9 (+8.8) |
4.2 Fine-grained Classification
To further investigate the generalizability of SeiT++, we conduct experiments on four different fine-grained datasets: Flowers [37], StandfordCars [29], iNat-18 [51], and iNat-19 [52]. Specifically, we tokenize each dataset and then fine-tune the token-trained model (ViT-B) on ImageNet-1k with the tokenized dataset. Since fine-grained datasets require the classification of objects with intricate characteristics, they are particularly sensitive to undesirable noise (i.e., RRC in SeiT) or excessive perturbation (i.e., Emb-Noise) resulting from data augmentation in the token embedding space. Table 3 shows the quantitative performance (top-1 accuracies) on fine-grained datasets. SeiT++ consistently outperforms the performance of SeiT across various datasets. These results confirm the effectiveness of our proposed method in fine-grained classification.
4.3 ADE-20k Semantic Segmentation
A limitation of token-based learning is often substantiated by learning dense prediction tasks. We conjecture that our method could facilitate training for those tasks, such as semantic segmentation. We extract tokens from ADE-20k [65] and use the pre-trained ViT-B/16 model on tokenized ImageNet-1k to evaluate the effectiveness of our method in semantic segmentation. We then fine-tune the pre-trained model using the tokenized ADE-20k dataset. Table 4 reports the mean intersection-over-union (mIoU) as an evaluation metric. The results show a significant performance improvement (+4.2%p mIoU) over SeiT. These results indicate that SeiT++ also significantly improves the model performance in the token-based dense prediction task.
@@ 4.3 ADE-20k Semantic Segmentation
| SeiT [38] | 74.0 | 50.7 | 62.6 | 25.5 | 22.6 | 15.7 |
| SeiT++ | 75.5 (+1.5) | 58.6 (+7.9) | 66.8 (+4.2) | 30.2 (+4.7) | 27.7 (+5.1) | 18.3 (+2.6) |
@@ 4.4 Evaluation on Robustness Datasets
To validate the effectiveness of our augmentation strategies in terms of robustness, we further evaluate the robustness of our method on a set of benchmarks such as adding Gaussian noise [25] or Gaussian blur [25], ImageNet-R [24], Sketch [53], and ObjectNet [3]. Notably, we adopt SeiT++ w/o MTM to validate the impact of our augmentation strategies on model robustness. The results in Table 5 show that SeiT++ consistently outperforms SeiT in terms of robustness accuracy. Previous studies [7, 49] observe that the key to the input pixel robustness depends on the pixel-level augmentations. Since our method enables pixel-based data augmentation to tokens, it achieves higher robustness accuracies than the baseline, which does not effectively exploit the pixel-based data augmentation.
4.5 Classification with Limited Data
Data augmentation becomes more critical in limited data scenarios. To assess the impact of our approach in such situations, we conduct experiments on the ImageNet-100 benchmark by randomly sampling data based on varying data ratios. The results (top-1 accuracies), as shown in Table 6, consistently demonstrate the superiority of SeiT++ over SeiT. Notably, the performance gap increases as the amount of training data decreases. With only 13k images, our method shows an improvement of approximately 21% over the baseline, indicating the effectiveness of our approach, especially in data-scarce scenarios.
@@ 4.5 Classification with Limited Data
| SeiT [38] | 77.3 | 70.3 | 53.3 | 43.2 |
| SeiT++ | 81.4 (+4.1) | 76.5 (+6.2) | 61.6 (+8.3) | 52.3 (+9.1) |
@@
4.6 Ablation Study
4.6.1 Impact of Each Augmentation Strategy.
In Table 7, we report the impact of proposed token augmentations. In particular, incorporating ColorAdapt yiqualds a 1.0%p performance improvement, highlighting the positive effect of preserving object structure during model training. Additionally, incorporating TokenAdapt yields an extra 3.1%p improvement. These results confirm that exploring token augmentations is a key factor in fully exploiting pre-extracted tokens.
| Input | SeiT | SeiT++ |
@@ 4.6 Ablation Study
4.6.2 Extensibility of Our Augmentation Strategy.
In this section, we conjecture the broad applicability of our augmentation strategies across various scenarios. To explore its extensibility, we delve into two distinct situations. Firstly, we examine the integration of our method with an alternative tokenizer—the OpenImages-trained VQGAN tokenizer [18], alongside the ImageNet-1k-trained ViT-VQGAN tokenizer [59]. The results in Table 8 consistently demonstrate improved performance on ImageNet-100 across different tokenizers. Additionally, we investigate the adaptability of our approach to a different input format by focusing on Discrete Cosine Transform (DCT) coefficients. Inspired by JPEG compression, we apply DCT to each image patch, yielding quantized DCT coefficients utilzed as input data during model training. As shown in Table 8, our approach enhances top-1 accuracy, similar to the tokenizer scenario. This consistent improvement across diverse input formats underscores the universal applicability of our method, highlighting its potential as a widely applicable solution for diverse input representations.
@@ 4.7 Implementation Details
We use the pre-trained ViT-VQGAN tokenizer [59] for token extraction. For our TokenAdapt module, we employ a single transformer block for conversion and reverse function. The TokenAdapt module is trained for one epoch on ImageNet-1k with a batch size of 128, a learning rate of 0.001 with cosine scheduling, and a weight decay of 0.05. We follow the training recipes from MAGE [32], SeiT [38], and mmsegmentation [9] for masked token modeling (MTM), token-based classification, and semantic segmentation, respectively. Detailed hyperparameters and settings are available in the Appendix.
@@ 5 Conclusion
In this paper, we investigate the compatibility of SeiT with self-supervised learning approaches in token-based frameworks. Among various self-supervised algorithms, Masked Image Modeling (MIM) [23, 58] has recently gained popularity for learning visual representations. Inspired by this, we extend SeiT by integrating Masked Token Modeling (MTM) for self-supervised pre-training. Additionally, recognizing that self-supervised approaches often require more data due to the lack of labels, we propose simple yet effective token-based augmentation strategies: TokenAdapt and ColorAdapt. TokenAdapt aligns the token embedding space with another embedding space compatible with pixel-based data augmentations. In addition, we introduce ColorAdapt, a color-changing augmentation for tokens. Our experimental results across various scenarios, including storage-efficient ImageNet-1k classification, fine-grained classification, ADE-20k semantic segmentation, and robustness benchmarks, consistently validate the effectiveness of our proposed method. Furthermore, the consistent performance improvement across diverse input formats highlights the potential of our approach as a widely applicable solution for various input representations.
Limitations. While SeiT++ is capable of handling large-scale data without labels with optimized storage requirements, experiments on such large datasets (e.g., ImageNet-21k [14]) are not contained in this paper due to resource constraints. Nevertheless, the advantages of MTM as self-supervised learning and the integration of our data augmentation strategies can be seamlessly integrated into larger datasets.
Supplementary Material
@@
In this supplementary material, we describe the details of SeiT++. Section 0.A examines the linear separability (Section 0.A.1), loss curve analysis (Section 0.A.2), data efficiency (Section 0.A.3), and computational cost (Section 0.A.4) of the proposed token augmentation strategy. Additionally, Section 0.B presents more qualitative results showing the effectiveness of our proposed TokenAdapt and ColorAdapt for augmenting tokens. Lastly, Section 0.C provides implementation details for our experiments.
Appendix 0.A Analysis of TokenAdapt
0.A.1 Linear Separability of TokenAdapt
To demonstrate the efficacy of our token augmentation strategies, we visualize the embedded feature representation in Figure A. This illustration offers insight into the linear separability of augmented data points, particularly when pixel-based data augmentation is applied through our proposed method. When pixel-based data augmentation is directly applied to token embedding without our method, there is significant overlap among augmented data points from different classes, indicating low linear separability. This overlap suggests that augmenting tokens without our method leads to substantial shifts in the distribution of training data. Maintaining distributional similarity between clean and augmented data is crucial for model performance, as witnessed in Cubuk et al. [13], and significant distribution shifts can result in performance degradation. In contrast, our method shows high linear separability, demonstrating its effectiveness in diversifying the data while minimizing the distributional gap between clean and augmented data. Moreover, this behavior aligns with the observed trend in the case of applying data augmentation to images (DeiT-B [50] in Figure A).
0.A.2 Loss Analysis of TokenAdapt
We analyze the influence of our approach on the training of ViT-B for both MTM pre-training and token-based ImageNet-1k classification, as illustrated by the loss curves in Figure B. Observing the training loss, our analysis shows that SeiT++, enhanced by our token augmentation strategy, maintains higher training loss values than SeiT [38] in both MTM pre-training and token-based ImageNet-1k classification. This pattern indicates that our method’s token augmentation enriches the data diversity, which helps to mitigate overfitting. Meanwhile, Figure B(c) shows that the test loss aligns with the trends of SeiT. Notably, while SeiT’s test loss increases in later training epochs, our approach consistently decreases, highlighting its effectiveness in preventing the overfitting problem.
| # of training data | 100 | 1k | 5k | 256k | 1281k |
@@ 0.A.2 Loss Analysis of TokenAdapt
0.A.3 Data Efficiency of TokenAdapt
To investigate the data efficiency of the TokenAdapt module, we explore the impact of reduced training data for the TokenAdapt module training. Specifically, we vary the number of training data for the TokenAdapt module training and subsequently train a token-based ImageNet-1k classification model using the trained TokenAdapt module. Table A shows that even with a reduced number of training data for our TokenAdapt module, the performance degradation is minimal. Notably, the top-1 accuracy remains more than 1%p higher than the baseline SeiT (74.0% top-1 accuracy), even with only 100 training data. This robust performance suggests that our TokenAdapt module is not significantly affected by variations in the training data size.
@@ 0.A.3 Data Efficiency of TokenAdapt
0.A.4 Computational Costs
@@
To validate the computational efficiency of TokenAdapt, we compare ViT-S models trained on the tokenized ImageNet-100 with different training strategies (Table B). Three scenarios are considered: (1) token-based augmentation only (i.e., SeiT), (2) pixel-based augmentations in decoded images (SeiT†), and (3) pixel-based augmentations using our TokenAdapt module. For SeiT†, using the ViT-VQGAN decoder (86M) and encoder (86M) during each forward computation significantly increases training time by over 8 times compared to the original SeiT training, making it impractical for addressing the data augmentation challenge. In contrast, our TokenAdapt module, requiring only 3.7M parameters, increases training time by only 8% compared to SeiT while achieving a remarkable performance improvement over 3%p in top-1 accuracy. Notably, TokenAdapt even outperforms SeiT†, indicating that the full decoding-augmentation-encoding process may introduce undesirable noise during tokenization. These results demonstrate the efficient handling of the data augmentation challenge by our TokenAdapt module. Notably, the TokenAdapt module can be removed during the inference.
Appendix 0.B More Qualitative Examples
0.B.0.1 TokenAdapt.